OpenStudy (anonymous):

Help is really needed For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (anonymous):

translate as "solve for \(x\), \( 3x+2=14\)

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

yep....x = 4 because 3(4) + 2 = 12 + 2 = 14

OpenStudy (anonymous):

so that would be the answer?

OpenStudy (texaschic101):

yes...when x = 4, then f(x) = 14

OpenStudy (anonymous):

ok can you help me with one more question please

OpenStudy (texaschic101):

I can try, but I am not the best at functions

OpenStudy (anonymous):

ok What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (texaschic101):

lets first find the slope of the line by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (3,6)...x1 = 3 and y1 = 6 (1,-2)..x2 = 1 and y2 = -2 now we sub slope(m) = (-2 - 6) / (1 - 3) slope(m) = -8/-2 slope(m) = 4 A parallel line will contain the same slope. Now use y = mx + b slope(m) = 4 (-1,1)...x = -1 and y = 1 now sub 1 = 4(-1) + b 1 = -4 + b 1 + 4 = b 5 = b so your parallel equation going through points (-1,1) is : y = 4x + 5

OpenStudy (anonymous):

So wait what's the slope of the line? y = 4x + 5

OpenStudy (texaschic101):

no...the slope of the line is 4. the parallel line containing points (-1,1) is : y = 4x + 5

OpenStudy (anonymous):

ok so the final answer is 4?

OpenStudy (texaschic101):

I used the slope formula , (y2 - y1) / (x2 - x1) to find the slope of the two points (3,6) and (1,-2). The reason that I had to find the slope is because I needed it to find the parallel line. The final answer is : y = 4x + 5, unless you need it in standard form

OpenStudy (anonymous):

No it doesn't need to be in standard. Thank you so much this really, really helped.

OpenStudy (texaschic101):

glad to help :)